# STELAR-VISION: Self-Topology-Aware Efficient Learning for Aligned Reasoning in Vision

### v2

@@ 4.5 Why Our Method Works?
Our method achieves substantial improvements on both in-distribution and out-of-distribution benchmarks. We hypothesize two key reasons for this success. First, for models that already possess the ability to generate diverse reasoning topologies but lack the capability to select the optimal topology for a given problem (e.g., Qwen2VL-7B-Instruct), our approach enables the model to adaptively choose the most effective structure. Second, for models that might not inherently support diverse topologies, our framework could instill this ability through guided supervision and reinforcement learning, making our methodology broadly applicable. However, due to computational constraints, we will leave systematic study to isolate this effect for future work.
@@ 5 Conclusion and Discussion
In this work, we propose STELAR-Vision, a training framework that enables VLM’s topology-aware reasoning ability via generated responses. STELAR-Vision enhances vision-language reasoning by leveraging diverse topological structures, achieving a 9.7% accuracy improvement over its base model and outperforming its larger variant Qwen2VL-72B-Instruct by 7.3%. The Frugal Learning reduces output length by 18.1% while maintaining comparable accuracy, surpassing Chain-Only baselines in both efficiency and task effectiveness. STELAR-Vision demonstrates strong generalization capabilities across OOD tasks, as evidenced by five diverse datasets spanning multiple domains. Compared to Chain-Only training, our method achieves 4.3% higher overall accuracy on in-distribution datasets and delivers consistent performance improvements across all OOD evaluations. These results highlight the benefits of topological augmentation and the model’s ability to adapt to a variety of reasoning challenges through dynamic reasoning structures.
Despite these promising results, STELAR-Vision relies on predefined topology types, and the dynamic relationship between problem structure and optimal reasoning topology remains underexplored. Future work will focus on enabling scalable, end-to-end topology induction and extending applicability to more advanced multimodal reasoning tasks.

### v3

@@ 4.5 Why Our Method Works?
Our method achieves substantial improvements on both in-distribution and out-of-distribution benchmarks. We hypothesize two key reasons for this success. First, for models that already possess the ability to generate diverse reasoning topologies but lack the capability to select the optimal topology for a given problem (e.g., Qwen2VL-7B-Instruct), our approach enables the model to adaptively choose the most effective structure. Second, for models that might not inherently support diverse topologies, our framework could instill this ability through guided supervision and reinforcement learning, and thus our methodology is still universal. However, due to computational constraints, we have not yet conducted a systematic study to isolate this effect, which we leave for future work.
@@ 5 Conclusion and Discussion
In this work, we propose STELAR-Vision, a training framework that enables topology-aware reasoning in VLMs via generated responses. STELAR-Vision enhances vision-language reasoning by leveraging diverse topological structures, achieving a 9.7% accuracy improvement over its base model and outperforming its larger variant Qwen2VL-72B-Instruct by 7.3%. The Frugal Learning variant reduces output length by 18.1% while maintaining comparable accuracy, surpassing Chain-Only baselines in both efficiency and task effectiveness. STELAR-Vision demonstrates strong generalization across five diverse OOD datasets and achieves 4.3% higher overall accuracy on in-distribution tasks, consistently outperforming Chain-Only training.
Despite these promising results, STELAR-Vision currently relies on predefined topology types, and the dynamic relationship between problem structure and optimal reasoning topology remains underexplored. Future work will focus on enabling scalable, end-to-end topology induction and extending applicability to more sophisticated multimodal reasoning tasks.
Supplementary Material
Example Prompts for TopoAug:
Figure: Prompts used in TopoAug to generate training samples for the MathVision (MATH-V) and VLM_S2H datasets. For each question, we prompt the generation models to produce 10 to 20 responses across three distinct reasoning topologies—Chain, Tree, and Graph—with extensive flexibility in maximum depth, number of children and neighbors.